# Expanding the Chaos: Neural Operator for Stochastic (Partial) Differential Equations

### v1

@@ Abstract
Stochastic differential equations (SDEs) and stochastic partial differential equations (SPDEs) are fundamental tools for modeling stochastic dynamics across the natural sciences and modern machine learning. Developing deep learning models for approximating their solution operators promises not only fast, practical solvers, but may also inspire models that resolve classical learning tasks from a new perspective. In this work, we build on classical Wiener–chaos expansions (WCE) to design neural operator (NO) architectures for SPDEs and SDEs: we project the driving noise paths onto orthonormal Wick–Hermite features and parameterize the resulting deterministic chaos coefficients with neural operators, so that full solution trajectories can be reconstructed from noise in a single forward pass. On the theoretical side, we investigate the classical WCE results for the class of multi-dimensional SDEs and semilinear SPDEs considered here by explicitly writing down the associated coupled ODE/PDE systems for their chaos coefficients, which makes the separation between stochastic forcing and deterministic dynamics fully explicit and directly motivates our model designs. On the empirical side, we validate our models on a diverse suite of problems: classical SPDE benchmarks, diffusion one-step sampling on images, topological interpolation on graphs, financial extrapolation, parameter estimation, and manifold SDEs for flood prediction, demonstrating competitive accuracy and broad applicability. Overall, our results indicate that WCE-based neural operators provide a practical and scalable way to learn SDE/SPDE solution operators across diverse domains.
@@ 1 Introduction
Stochastic Differential Equations (SDEs) and Stochastic Partial Differential Equations (SPDEs) are fundamental tools for modelling complex dynamical systems across different scientific domains. Examples range from the Ginzburg–Landau equation for superconductivity (temam2012infinite) and the stochastic Navier–Stokes equations for incompressible fluid dynamics in SPDEs (boyer2012mathematical) to the Ornstein–Uhlenbeck (OU) process for modelling the velocity of Brownian particles in statistical mechanics (martin2021statistical) and the Heston model for financial volatility forecasting (de2018machine).
@@
Recently, these equations have emerged as cornerstones in machine learning. A prominent example is the OU process, whose time-reversal properties now form the theoretical backbone of diffusion generative models, which have achieved remarkable results in image generation (song2020score), data interpolation (de2021diffusion), and time series forecasting (lin2024diffusion). However, despite their broad applicability, developing deep learning tools for approximating solutions to these equations, or a fundamental understanding of their complex solution structures, remains in its early stages (neufeld2024solving).
@@
With the recent success of NOs for learning mappings between function spaces (li2020fourier; kovachki2023neural), a growing literature has applied NOs to learn the solution operators of SPDEs and SDEs, including Neural SPDE (salvi2022neural), GINO (li2023geometry), FNO (li2020fourier), and DeepONet (lu2019deeponet). Their key advantage is one-shot evaluation: solutions at any query time can be obtained from a single forward pass, rather than iterative time stepping. However, incorporating stochasticity remains non-trivial (salvi2022neural). The general strategy is to condition on the stochastic forcing terms by providing the noise path or its features, thereby reducing the problem to learning a deterministic operator.
@@
Building on this principle, we propose novel NO models for SPDEs and SDEs grounded in the Wiener Chaos Expansion (WCE). By projecting the driving noise onto orthonormal Wick–Hermite features and conditioning on them, our models learn deterministic operators that disentangle stochastic forcing from deterministic dynamics, thereby reconstructing full trajectories. On the theoretical side, we build on classical WCE representations for the class of multi-dimensional SDEs and semilinear SPDEs by explicitly writing down the coupled ODE/PDE systems for their chaos coefficients. In particular, we show how the propagator system takes an explicit ODE form for SDEs driven by multi-dimensional Brownian motion and a corresponding PDE form for SPDEs. We leverage these formulations directly to motivate our NO parameterization. Empirically, we evaluate our models on classical SPDE benchmarks and several SDE-based downstream tasks (diffusion one-step sampling, graph interpolation, financial extrapolation, parameter estimation, and manifold SDEs), achieving competitive accuracy and strong agreement with reference solution operators across all settings.
@@
We develop a NO framework for SPDEs and SDEs that is built on classical Wiener–chaos expansions: solutions are represented via chaos coefficients, and the corresponding deterministic propagator functions are parameterized by neural operators. On the analytical side, we restate and slightly extend known WCE representations by deriving explicit coupled ODE/PDE systems for their Wiener–chaos coefficients. On the empirical side, we apply this framework to classical SPDE benchmarks and many SDE-based downstream tasks, and observe competitive accuracy with one-shot evaluation. To the best of our knowledge, this is the first work to systematically deploy NO models for both SPDE and SDE solution operators across such a diverse and extensive set of experiments.
@@ 2 Mathematical Preliminaries
This section introduces the preliminaries for SPDE and SDE, following the standard notation in (neufeld2024solving; huschto2014solving; eigel2024functional).
@@
Given $T>0$ and a filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$, we consider the following semi-linear SPDE [^1] An SPDE is linear if both $A,F,B$ are linear, or semi-linear if $A$ is linear but $F,B$ are non linear.:
$dX_{t}=(AX_{t}+F(t,\cdot,X_{t}))dt+B(t,\cdot,X_{t})dW_{t},\quad X_{0}=\chi_{0}\in\mathcal{H},$ (SPDE)
@@
$X_{t}=S_{t}\chi_{0}+\int_{0}^{t}S_{t-s}F(s,\cdot,X_{s})ds+\int_{0}^{t}S_{t-s}B(s,\cdot,X_{s})dW_{s},$ (1)
holds almost surely. Here $S=e^{tA}:\mathcal{H}\to\mathcal{H}$ refers to a $C_{0}$-semigroup generated by operator $A$. In addition, we will denote a single realization of the solution process for a fixed random event $\omega\in\Omega$ by $u(t,x)$, where $x\in\mathcal{D}$ is the spatial variable.
@@
The SDE is a special form of SPDE with $A=0$ and $\mathcal{H}$, $\widetilde{\mathcal{H}}$ reduced to finite-dimensional vector spaces. Specifically,
$dX_{t}=F(t,X_{t})dt+B(t,X_{t})dW_{t},\quad X_{0}=x_{0}\in\mathbb{R}^{d},$ (SDE)
where $W_{t}$ is an $m$-dimensional Brownian motion with covariance $Q\in\mathbb{R}^{m\times m}$ (i.e., $\mathbb{E}[(W_{t_{2}}-W_{t_{1}})(W_{t_{2}}-W_{t_{1}})^{\top}]=(t_{2}-t_{1})Q$), and $F(t,\cdot)\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$, $B(t,\cdot):\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are known as the drift and diffusion term of the SDE, respectively. Without loss of generality, in this work, we only consider the standard Brownian motion, i.e., $Q=I$, yet our conclusions can be naturally extended to general cases. The strong solution of (SDE) is
$X_{t}=x_{0}+\int_{0}^{t}F(s,X_{s})ds+\int_{0}^{t}B(s,X_{s})dW_{s}.$ (2)
Analogous to SPDE, we denote a single realization of the solution process $X_{t}$ by the deterministic vector-valued function $u(t)$.
@@
3 Wiener Chaos Expansion of SPDEs and SDEs
In this section, we recall the WCE for SDEs and SPDEs and adapt it to the setting of this paper. We first review how Brownian and Q-Brownian motions can be represented via countable Gaussian coordinates constructed from Brownian increments, and then show how this representation leads to WCEs of SDE/SPDE solutions. Our presentation follows the classical theory (luo2006wiener); we restate and slightly refine the main ingredients needed for our NO construction, in particular by making the associated propagator systems explicit, all proofs are in Appendix C for completeness.
@@ 3.1 Reconstruction of Brownian Motions
We begin by illustrating some standard conclusions on reconstruction of Brownian and Q-Brownian motions through their increments (da2014stochastic).
Figure 1: Reconstruction of a one-dimensional Brownian motion and a two-dimensional Q-Brownian motion from truncated chaos expansions with different truncation orders $n$. The blue curve shows the true trajectory, while the coloured curves show reconstructions using $n$ temporal modes. For very small $n$ (e.g., $n=5$) the approximation is intentionally coarse and appears oversmoothed, especially in the second coordinate of the Q-Brownian motion, but for larger $n$ the reconstructions closely track the true paths, consistent with Lemmas 1 and 2.
@@
Let $\{e_{j}\}_{j\in\mathbb{N}}\subset L^{2}([0,T])$ be a complete orthonormal family that is continuous almost everywhere. Let $W=(W^{(1)},\dots,W^{(d)})$ be a $d$-dimensional standard Brownian motion on $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. For every $i\in\{1,\dots,d\}$ and $j\in\mathbb{N}$, define $\xi_{ij}\coloneqq\int_{0}^{T}e_{j}(s)\,dW^{(i)}_{s}$. Then the collection $\{\xi_{ij}\}_{i,j}$ consists of mutually independent standard Gaussian random variables (neufeld2024solving).
@@
We now extend our conclusion to show how the Q-Brownian motion in $\mathcal{H}$ can be reconstructed.
@@ 3.2 Chaos Expansion on SPDEs and SDEs
We now recall how the SPDEs, SDEs solutions can be coupled with a polynomial basis constructed from the Gaussian random variables introduced above. We start by defining the Wick polynomials.
Definition 2 (Wick Polynomial).
Let $\Xi=\{\xi_{mj}\}_{m,j\in\mathbb{N}}$ be the family of Gaussian random variables constructed from Definition 1. The index $m$ corresponds to the spatial/component mode, while $j$ relates to the temporal basis. Let $\mathcal{J}$ be the set of all multi-indices $\alpha=(\alpha_{mj})_{m,j\in\mathbb{N}}$ with finite support. For any $\alpha\in\mathcal{J}$, the corresponding normalized Wick monomial is
@@
The family $\{\xi_{\alpha}\}_{\alpha\in\mathcal{J}}$ forms a complete orthonormal basis for the space of square-integrable random variables measurable with respect to $\Xi$, i.e., $\mathbb{E}[\xi_{\alpha}\xi_{\beta}]=\delta_{\alpha\beta},\,\,\forall\alpha,\beta\in\mathcal{J}$. Furthermore, the linear span of this basis is dense in $L^{p}(\Omega,\mathcal{F},\mathbb{P})$ for any $p\in[1,\infty)$ (neufeld2024solving). This completeness is fundamental, as it allows any random variable in this space to be represented as a series expansion in this basis. This representation effectively transforms stochastic equations into an infinite system of deterministic equations for the coefficients of the expansion. We first apply this framework to the case of SDEs before generalizing to the more complex setting of SPDEs.
Theorem 1 (SDE Propagator System).
Let $X_{t}\in\mathbb{R}^{d}$ be the strong solution to the $d$-dimensional SDE (SDE) driven by an $m$-dimensional Brownian motion $W_{t}$. The solution’s WCE is given by
@@
The one-dimensional version of this propagator system for (SDE) has been used in the SDEONet framework of (eigel2024functional). Building on classical WCE results, here we write out the multi-dimensional Brownian case explicitly: the chaos indices now carry both a component index $k$ and a temporal basis index $i$, and the right-hand side of Eq. (5) depends only on $t$ (and on fixed model coefficients), not on the randomness $\omega$. Each coefficient $u_{\alpha}(t)$ therefore solves a deterministic ODE in time, which makes these propagators natural targets for neural networks that take $t$ as input. The same idea extends to SPDEs: Theorem 2 shows that the chaos coefficients become deterministic space–time fields $u_{\alpha}(t,x)$ satisfying a coupled system of PDEs.
@@
The evolution of propagator fields $u_{\alpha}(t,x)$ for (SPDE) is governed by a coupled system of deterministic PDEs
@@
As discussed in the SPDE WCE literature (see, e.g., Remark 2.13 in neufeld2024solving), and (lototsky2006stochastic) the chaos coefficients of semilinear SPDEs satisfy a coupled system of deterministic evolution equations. Theorem 2 makes this structure explicit for the class of SPDEs in (SPDE) by writing the propagator system in the form (6); examples and concrete expressions for $\mathscr{F}_{\alpha}$ are given in Appendix C. While the precise form of $\mathscr{F}_{\alpha}$ depends on the choice of $F$ and $B$, the key point for our purposes is that each $u_{\alpha}(t,x)$ solves a deterministic PDE in space–time. This observation enables us to use standard deterministic PDE neural operators such as FNO (li2020fourier) and GINO (li2023geometry) to approximate the propagator fields, with the stochasticity handled entirely by the Wick features of the noise trajectories.
Figure 2: General working flow of $\mathcal{F}$-SPDENO: a discrete Q-Brownian path is simulated, Wick features computed from its increments are concatenated with the initial condition $\chi_{0}$ and passed through an FNO to produce time-independent propagator fields, which are combined with a fixed temporal basis to reconstruct the full trajectory.
4 Model Architecture
$\mathcal{F}$-SPDENO As discussed above, solutions of (SPDE) admit a Wiener–chaos representation $X_{t}(x)=\sum_{\alpha\in\mathcal{J}}u_{\alpha}(t,x)\,\xi_{\alpha}$, so that once the Q-Brownian motion $W$ and the initial condition $\chi_{0}$ are observed, the randomness is entirely carried by the chaos variables $\{\xi_{\alpha}\}$, while the propagators $u_{\alpha}(t,x)$ remain deterministic. In practice, we simulate a discrete Q-Brownian path on a time grid, compute its increments, and from them construct the Gaussian variables $\{\xi_{mj}\}$ and Wick monomials $\{\xi_{\alpha}\}$; a truncated collection of $\{\xi_{\alpha}(W)\}$ is stacked into the Wick feature vector $\eta(W)$ (Section 2), which serves as the stochastic input to our model. To make the propagators suitable NO targets, we further expand them in a temporal basis $\{\phi_{k}(t)\}_{k=1}^{K}$ as $u_{\alpha}(t,x)\approx\sum_{k=1}^{K}u_{\alpha,k}(x)\,\phi_{k}(t)$, so that we only need to learn the time-independent coefficient fields $\{u_{\alpha,k}(x)\}$. $\mathcal{F}$-SPDENO then uses an FNO with parameters $\theta$ to approximate the deterministic operator $(\chi_{0},\eta(W))\mapsto\{u_{\alpha,k}(x)\}$, and the final prediction is obtained by combining the FNO outputs with the chaos and temporal bases. Figure 2 summarises this workflow: the block labelled “Deterministic PDE” is a conceptual illustration of the propagator system in Theorem 2, but in practice we do not solve this PDE system explicitly and instead train the FNO to approximate the mapping $(\chi_{0},\eta(W))\mapsto\{u_{\alpha,k}(x)\}$ directly.
SDENO For the SDE case, we follow the same chaos-based representation and treat the propagators as deterministic functions of time that can be approximated by standard neural networks. As illustrated in the previous section, once the Brownian motion has been encoded into Wick features, each chaos coefficient $u_{\alpha}(t)$ satisfies a deterministic ODE in $t$, so one can use a wide range of architectures (even simple MLPs) to approximate these time-dependent propagators. Concretely, we construct Wick features from the driving noise trajectory $\{W_{t}\}_{t\in[0,T]}$ and feed the time variable $t$ (or its positional encoding) into a backbone network, whose outputs are then combined with the Wick features via an explicit inner product. This yields a flexible SDENO family with different backbones: for example, we use a UNet for diffusion SDEs on images ($\mathcal{U}$-SDENO) and graph neural networks for diffusion interpolation on graphs ($\mathcal{G}$-SDENO). In $\mathcal{U}$-SDENO, the time dependence is modeled inside the UNet, whereas in $\mathcal{G}$-SDENO it is handled by a separate temporal network. Further architectural details and additional experiments for the SDE settings are provided in the appendix.
5 Numerical Experiments
In this section, we present a comprehensive empirical evaluation of our neural operator (NO) models across five settings: (i) solving SPDEs (the dynamic $\boldsymbol{\Phi}^{4}_{1}$ model and the 2D Navier-Stokes equation); (ii) solving SDEs in the image and graph domains, yielding a diffusion one step sampler and a topological interpolator; (iii) extrapolation on the financial Heston model; (iv) parameter estimation with our generalized Meta-SDENO model; (v) learning SDE solution on manifold. All experiments are executed on an NVIDIA® H200 SXM GPU (141 GB HBM3e) within an HPC cluster. Additional experimental details and results, e.g., empirical parameter sensitivity, are provided in Appendix F.
5.1 Approximation of SPDE Solutions
Dynamic $\boldsymbol{\Phi}^{4}_{1}$ Model
We begin to evaluate our $\mathcal{F}$-SPDENO model by the dynamic $\boldsymbol{\Phi}^{4}_{1}$ model, which is widely applied in the realm of superconductivity (temam2012infinite). The equation has the form
$dX_{t}=\bigl(\Delta X_{t}+3X_{t}-X_{t}^{3}\bigr)\,dt+dW_{t},\,\,X(0,x)=\chi_{0}\in L^{2}(\mathbb{T}^{1}),\quad X_{t}(0)=X_{t}(1),$ (7)
where $(t,x)\in[0,T]\times[0,1]$, and $W$ is the Q-Brownian motion. We follow the experimental settings of salvi2022neural; hu2022neural in generating sample paths, initial conditions, ground truth solutions, and train-test splits. We conduct two tasks with different training observations $N=1000$ and $10000$, respectively, and evaluate our model with two settings, (i) $W\mapsto X$ which assumes $W$ is observed but the initial condition is fixed all the time and (ii) $(X_{0},W)\mapsto X$ where $W$ is observed and $X_{0}$ changed across the samples. We compare $\mathcal{F}$-SPDENO with various baselines such as NCDE, NCDE-FNO, NRDE (lu2022comprehensive), FNO (li2020fourier) and NSPDE (salvi2022neural). We note that we did not compare our model to architectures specifically designed for rough or even singular SPDEs, such as (hu2022neural; gong2023deep). In the one-dimensional dynamic $\boldsymbol{\Phi}^{4}_{1}$ setting considered here, the solution enjoys positive Hölder regularity (li2025spdebench), so the regularity-structure machinery that underpins these models is not essential. The results are contained in Table 1. Our model achieves remarkable results relative to all other baselines, with lower $L^{2}$ errors across both small and large training schemes. Interestingly, increasing $N$ from $1000$ to $10000$ only leads to modest changes in the relative $L^{2}$ error for NSPDE and $\mathcal{F}$-SPDENO, especially on the $(X_{0},W)\mapsto X$ task. This behaviour is consistent with the fact that both methods already achieve very small errors in the $N=1000$ regime. In $\boldsymbol{\Phi}^{4}_{1}$, $N=1000$ trajectories already capture most of the variability of the dynamics, so increasing $N$ to $10000$ brings only limited additional benefit. This aligns with the recent benchmarking study in SPDE (li2025spdebench).
@@ 5.1 Approximation of SPDE Solutions
$N=1000$ $N=10000$
Model #Parameters Inference time (s) $W\mapsto X$ $(X_{0},W)\mapsto X$ $W\mapsto X$ $(X_{0},W)\mapsto X$
NCDE 272 672 0.503 0.134 0.168 0.050 0.064
NRDE 2 164 356 0.105 0.129 0.150 0.058 0.089
NCDE-FNO 48 769 1.845 0.077 0.061 0.091 0.060
FNO 1 647 105 0.165 0.029 0.034
NSPDE 265 089 0.248 0.021 0.020 0.013 0.021
$\mathcal{F}$-SPDENO  (Ours) 108 629 0.128 0.012 0.015 0.011 0.017
Navier-Stokes Equation
In the second experiment, we evaluate $\mathcal{F}$-SPDENO on the 2D incompressible Navier-Stokes equations in vorticity form (li2020fourier). The dynamic takes the form
Here we take $U[X_{t}]\coloneqq\nabla^{\perp}(-\Delta)^{-1}X_{t}$ with $\nabla^{\perp}\coloneqq(-\partial_{x_{2}},\partial_{x_{1}})$ where $\partial$ denotes the partial derivative of the spatial variable (i.e., $x=(x_{1},x_{2})$). The parameter $\nu>0$ is the viscosity; $f$ is a deterministic vorticity forcing. Similarly, we generate the samples and ground truth solutions, followed by the work in (salvi2022neural). We train and evaluate our model on $\times64$ grid.

### v2

@@ Abstract
Stochastic differential equations (SDEs) and stochastic partial differential equations (SPDEs) are fundamental for modeling stochastic dynamics across the natural sciences and modern machine learning. Learning their solution operators with deep learning models promises fast solvers and new perspectives on classical learning tasks. In this work, We build on Wiener–chaos expansions (WCE) to design neural operator (NO) architectures for SDEs and SPDEs: we project driving noise paths onto orthonormal Wick–Hermite features and use NOs to parameterize the resulting chaos coefficients, enabling reconstruction of full trajectories from noise in a single forward pass. We also make the underlying WCE structure explicit for multi-dimensional SDEs and semilinear SPDEs by showing the coupled deterministic ODE/PDE systems governing these coefficients. Empirically, we achieve competitive accuracy across several tasks, including standard SPDE benchmarks and SDE-based diffusion one-step image sampling, topological graph interpolation, financial extrapolation, parameter estimation, and manifold SDE flood forecasting. These results suggest WCE-based neural operators are a practical and scalable approach to learning SDE/SPDE solution operators across domains.
Machine Learning, ICML
@@ 1 Introduction
Stochastic Differential Equations (SDEs) and Stochastic Partial Differential Equations (SPDEs) are fundamental tools for modelling complex dynamical systems across different scientific domains. Examples range from the Ginzburg–Landau equation for superconductivity (Temam, 2012) and the stochastic Navier–Stokes equations for incompressible fluid dynamics in SPDEs (Boyer and Fabrie, 2012) to the Ornstein–Uhlenbeck (OU) process for modelling the velocity of Brownian particles in statistical mechanics (Martin et al., 2021) and the Heston model for financial volatility forecasting (De Spiegeleer et al., 2018).
@@
Recently, these equations have emerged as cornerstones in machine learning. A prominent example is the OU process, whose time-reversal properties now form the theoretical backbone of diffusion generative models, which have achieved remarkable results in image generation (Song et al., 2020b), data interpolation (De Bortoli et al., 2021), and time series forecasting (Lin et al., 2024). However, despite their broad applicability, developing deep learning tools for approximating solutions to these equations, or a systematic understanding of their complex solution structures, remains in its early stages (Neufeld and Schmocker, 2024).
@@
With the recent success of NOs for learning mappings between function spaces (Li et al., 2020; Kovachki et al., 2023), a growing literature has applied NOs to learn the solution operators of SPDEs and SDEs, including Neural SPDE (Salvi et al., 2022), GINO (Li et al., 2023), FNO (Li et al., 2020), and DeepONet (Lu et al., 2019). Their key advantage is one-shot evaluation: solutions at any query time can be obtained from a single forward pass, rather than iterative time stepping. However, incorporating stochasticity remains non-trivial (Salvi et al., 2022). The general strategy is to condition on the stochastic forcing terms by providing the noise path or its features, thereby reducing the problem to learning a deterministic operator.
@@
Building on this principle, we propose NO models for SPDEs and SDEs grounded in the Wiener Chaos Expansion (WCE). By projecting the driving noise onto orthonormal Wick–Hermite features and conditioning on them, our models learn deterministic operators that disentangle stochastic forcing from deterministic dynamics, thereby reconstructing full trajectories. We make classical Wiener–chaos representations explicit by expressing the propagator dynamics as coupled deterministic ODE systems for multi-dimensional SDEs and PDE systems for semilinear SPDEs, which we use directly to guide our neural operator design. Empirically, we evaluate our models on classical SPDE benchmarks and several SDE-based downstream tasks (diffusion one-step sampling, topological interpolation, financial extrapolation, parameter estimation, and manifold SDEs), achieving competitive accuracy and strong agreement with reference solution operators across all settings.
@@
We develop a NO framework for SPDEs and SDEs that is built on classical Wiener–chaos expansions: solutions are represented via chaos coefficients, and the corresponding deterministic propagator functions are parameterized by neural operators. We explicitly formulate classical Wiener–chaos representations by writing out the coupled deterministic ODE/PDE systems for chaos coefficients (Theorem 1 and 2), which can be directly parameterized and learned by neural operators. We apply this framework to classical SPDE benchmarks and many SDE-based downstream tasks, and observe competitive accuracy with one-shot evaluation. To the best of our knowledge, this is the first work to systematically deploy NO models for both SPDE and SDE solution operators across such a diverse and extensive set of experiments.
@@ 2 Mathematical Preliminaries
We start by introducing the preliminaries for SPDEs. Given $T>0$ and a filtered probability space $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$, we consider the following semi-linear SPDE [^1] An SPDE is linear if both $A,F,B$ are linear, or semi-linear if $A$ is linear but $F,B$ are non linear.:
$dX_{t}=(AX_{t}+F(t,\cdot,X_{t}))dt+B(t,\cdot,X_{t})dW_{t},$ (1)
@@
The SDE is a special form of SPDE with $A=0$ and $\mathcal{H}$, $\widetilde{\mathcal{H}}$ reduced to finite-dimensional vector spaces.
$dX_{t}=F(t,X_{t})dt+B(t,X_{t})dW_{t},$ (2)
with initial condition $X_{0}=x_{0}\in\mathbb{R}^{d}$, and $W_{t}$ is an $m$-dimensional Brownian motion with covariance $Q\in\mathbb{R}^{m\times m}$ (i.e., $\mathbb{E}[(W_{t_{2}}-W_{t_{1}})(W_{t_{2}}-W_{t_{1}})^{\top}]=(t_{2}-t_{1})Q$), and $F(t,\cdot)\colon\mathbb{R}^{d}\to\mathbb{R}^{d}$, $B(t,\cdot):\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ are known as the drift and diffusion terms, respectively. Without loss of generality, in this work, we only consider the standard Brownian motion, i.e., $Q=I$, yet our conclusions can be naturally extended to general cases. The strong solution of Eq. (2) is $X_{t}=x_{0}+\int_{0}^{t}F(s,X_{s})ds+\int_{0}^{t}B(s,X_{s})dW_{s}$. Analogous to SPDE, we denote a single realization of the solution process $X_{t}$ by the deterministic vector-valued function $u(t)$.
@@
Figure 1: Reconstruction of a one-dimensional Brownian motion and a two-dimensional Q-Brownian motion by the truncation orders $n$.
3 WCE of SPDEs and SDEs
In this section, we summarize the classical Wiener Chaos Expansion (WCE) framework for SDEs and SPDEs with a focus on making it operational for neural operator learning. We describe how Brownian motion and Q-Brownian motion can be represented via countable Gaussian coordinates constructed from Brownian increments, and how this representation yields WCE-based solution forms for SDEs and SPDEs. We then present the corresponding deterministic propagator systems for the chaos coefficients explicitly, which serve as the targets learned by our neural operator models. All formal statements are collected in Appendix C.
@@ 3.1 Reconstruction of Brownian Motions
We begin by illustrating some standard conclusions on reconstruction of Brownian and Q-Brownian motions through their increments (Da Prato and Zabczyk, 2014; Neufeld and Schmocker, 2024).
@@
Let $\{e_{j}\}_{j\in\mathbb{N}}\subset L^{2}([0,T])$ be a complete orthonormal family that is continuous almost everywhere. Let $W=(W^{(1)},\dots,W^{(d)})$ be a $d$-dimensional standard Brownian motion on $(\Omega,\mathcal{F},\mathbb{F},\mathbb{P})$. For every $i\in\{1,\dots,d\}$ and $j\in\mathbb{N}$, define $\xi_{ij}\coloneqq\int_{0}^{T}e_{j}(s)\,dW^{(i)}_{s}$. Then the collection $\{\xi_{ij}\}_{i,j}$ consists of mutually independent standard Gaussian random variables (Neufeld and Schmocker, 2024).
@@
Similar conclusions can be built for Q-Brownian motions.
@@
Figure 2: General working flow of $\mathcal{F}$-SPDENO: Discretized Q-Brownian motion as the initial input, Wick features computed from its increments are concatenated with the initial condition $\chi_{0}$ and passed through an FNO to produce time-independent propagator fields, which are combined with a fixed temporal basis to reconstruct the full trajectory.
@@ 3.2 Chaos Expansion on SPDEs and SDEs
We now show how the SPDE and SDE solutions can be coupled with a polynomial basis constructed from the Gaussian random variables introduced above. We begin by introducing the Wick polynomials.
Definition 2.
Let $\Xi=\{\xi_{mj}\}_{m,j\in\mathbb{N}}$ be the family of Gaussian random variables constructed from Definition 1. Further let $\mathcal{J}$ be the set of all multi-indices $\alpha=(\alpha_{mj})_{m,j\in\mathbb{N}}$ with finite support. For any $\alpha\in\mathcal{J}$, the corresponding normalized Wick monomial is
@@
The family $\{\xi_{\alpha}\}_{\alpha\in\mathcal{J}}$ forms a complete orthonormal basis for the space of square-integrable random variables measurable with respect to $\Xi$, i.e., $\mathbb{E}[\xi_{\alpha}\xi_{\beta}]=\delta_{\alpha\beta},\,\,\forall\alpha,\beta\in\mathcal{J}$. Furthermore, the linear span of this basis is dense in $L^{p}(\Omega,\mathcal{F},\mathbb{P})$ for any $p\in[1,\infty)$ (Neufeld and Schmocker, 2024). This completeness is fundamental, as it allows any random variable in this space to be represented as a series expansion in this basis. Finally, we show the well-known WCE for SDE as follows.
Theorem 1 (WCE for multi-dimensional SDE).
Let $X_{t}\in\mathbb{R}^{d}$ be the strong solution to the $d$-dimensional SDE (2) driven by an $m$-dimensional Brownian motion $W_{t}$. The solution’s WCE is given by
@@
where $u_{\alpha}(t)\in\mathbb{R}^{d}$ are called the propagators of SDE, and $u(t)$ is governed by a coupled system of ODEs. That is
$\frac{d}{dt}u_{\alpha}^{(j)}(t)=$ $\mathbb{E}\left[F^{(j)}(t,X_{t})\xi_{\alpha}\right]$
$+$ $\sum_{k=1}^{m}\sum_{i=1}^{\infty}\sqrt{\alpha_{ki}}\,e_{i}(t)\,\mathbb{E}\left[B^{(j,k)}(t,X_{t})\xi_{\alpha-e_{ki}}\right],$ (5)
@@
The one-dimensional version of this propagator system for Eq. (2) has been studied in the SDEONet framework of (Eigel and Miranda, 2024). We show the form of ODEs that govern the propagators in the multi-dimensional case in Theorem 1. Since these ODEs depend only on time $t$, the corresponding propagators can be efficiently learned by neural networks (or ODE solvers) with $t$ as input, enabling direct reconstruction of the SDE solution. This idea extends to SPDEs as follows.
@@
The evolution of propagator fields $u_{\alpha}(t,x)$ for Eq. (1) is governed by a coupled system of deterministic PDEs
@@
Aligned with discussed in the SPDE WCE literature (e.g., Remark 2.13 in Neufeld and Schmocker, 2024), Theorem 2 makes the WCE structure for SPDE explicit in Eq. (1) by writing the propagator system in the form (6). We show some examples of $\mathscr{F}_{\alpha}$ in Appendix C. While the precise form of $\mathscr{F}_{\alpha}$ depends on the choice of $F$ and $B$, the key point for our purposes is that each $u_{\alpha}(t,x)$ solves a deterministic PDE in space–time. This observation enables us to use standard deterministic PDE neural operators such as FNO (Li et al., 2020) and GINO (Li et al., 2023) to approximate the propagator fields, with the stochasticity handled entirely by the Wick features (i.e., $\xi_{\alpha}$) of the noise trajectories.
@@
$N=1000$ $N=10000$ Model #Parameters Inference time (s) $W\mapsto X$ $(X_{0},W)\mapsto X$ $W\mapsto X$ $(X_{0},W)\mapsto X$ NCDE 272 672 0.503 0.134 0.168 0.050 0.064 NRDE 2 164 356 0.105 0.129 0.150 0.058 0.089 NCDE-FNO 48 769 1.845 0.077 0.061 0.091 0.060 FNO 1 647 105 0.165 0.029 0.034 NSPDE 265 089 0.248 0.021 0.020 0.013 0.021 $\mathcal{F}$-SPDENO (Ours) 108 629 0.128 0.012 0.015 0.011 0.017